Monthly spending ($)

 

 

 

 

 

 

Age

20s

30s

$0.00 - $49.99

75

29

$50.00 - $99.99

91

83

$100.00 - $149.99

113

90

$150.00 - $199.99

89

128

$200.00 - $249.99

63

159

$250.00 or more

34

46

Total customers

465

535

 

a. Given that a customer spends $200.00 or more, what is the probability that he/she is in his/her 30s?



1
Expert's answer
2021-05-13T01:14:58-0400

The given data is-





(a) Total number of customers who spends $ 200 and more-

           =63+34+159+46=302


 Number of customers who spends $ 200 and more in 30s =159+46=205


Required Probability "=\dfrac{205}{302}=0.679"